a fountain thatis 5ft tall sprays water into the air with an upward velocity of 22ft/s. What function gives the hieght of the water in feet and seconds after it is sprayed upward? what is the maximum height ofthe water?

Answer by stanbon(75887)

a fountain that is 5ft tall sprays water into the air with an upward velocity of 22ft/s. What function gives the height of the water in feet and seconds after it is sprayed upward? what is the maximum height ofthe water?
---------------
Equation:
h(t) = -16t^2+22t+5
-------------
Max occurs when t = -b/2a = -22/(2*-16) = 11/16 seconds
----
Max height = h(11/16) = 12.563 ft.
===========================
Cheers,
Stan H.